Etch Group morning equipment status: 12/12/2005
eenriquez at snf.stanford.edu
Tue Dec 13 00:00:21 PST 2005
* Etch Group - Morning Equipment Status Report
*12/13/2005 12:01 am*
*Etch Equipment :*
System still does not boot up. Currently suspect VME controller and
/or hard drive + control.
What has been done to date.
1) Repaired cause of water leak.
2) Wet cleaned the chamber.
3) Replaced damaged DC power supply.
4) Attempted to start-up the system with spare VME board ,hard drive
and controller (parts
condition was unknown).
5) Ordered new VME board. Should arrive 12/2.
-- VME did not fix the problem.
Field service believes the CPU board is bad. Ordered CPU board.
Bad dc power supplies. They should be in by Wed.
Stillll waiting for the final permit. Running tool in "Test" mode
More information about the etch